# Bregman Stochastic Proximal Point Algorithm with Variance Reduction

### v1

@@ B.2 Technical lemmas
Suppose that Assumptions 4.5 are verified and that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1 with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence. Then, for all $k\in\mathbb{N}$,
@@
Proof.
Define $\mathsf{E}_{k}[\cdot]=\mathsf{E}[\cdot\,|\,\mathfrak{F}_{k}]$. It follows from Proposition B.1 that
@@
$=\frac{1}{2}\Big[D_{h^{*}}\left(\nabla({\mathsf{x}}_{1}^{+}),\nabla h({\mathsf{x}})\right)+D_{h^{*}}\left(\nabla({\mathsf{x}}_{2}^{+}),\nabla h({\mathsf{x}})\right)\Big].$
@@
Lemma B.7 (Bregman variance decomposition by [24]).
Let x{x} be a random variable on $\mathbb{R}^{d}$. Then for any ${\mathsf{u}}\in\mathbb{R}^{d}$,
@@
B.3 Proofs of Section 4.1
Proof of Theorem 4.6. From Lemma B.4 and since $B_{k}-\rho M_{k}\leq 0$ , we get
@@
Proof of Theorem 4.7. From Lemma B.4 and relative strong convexity of $F$, we have
@@
B.4 Proofs of Section 5
Proof of Theorem 5.2. From Proposition B.1, we have
@@ B.4 Proofs of Section 5
Proof of Theorem 5.4. Proposition B.1 gives
@@
Proof of Lemma 5.8. From Equation (4.2), we infer that $\nabla h({z}_{k+1})=\nabla h({x}_{k})-\alpha_{k}\nabla f_{i_{k}}({x}_{k})+\alpha_{k}{e}_{k}$. Set $\nabla h({\mathsf{x}}_{1}^{+})=\nabla h({x}_{k})-2\alpha_{k}[\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*})]$ and $\nabla h({\mathsf{x}}_{2}^{+})=\nabla h({x}_{k})-2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})$. Then, from Lemma B.6, we know that $D_{h}({x}_{k},{z}_{k+1})\leq(S_{1}+S_{2})/2$, where
@@
Using the gain function in Equation (2.1), then Lemma B.5, and finally Equation (5.2), we have
@@
$\leq 4L\alpha_{k}^{2}\mathsf{E}_{k}\left[G\left(\nabla h({x}_{k}),\nabla h({x}_{k}),\frac{1}{L}(\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*}))\right)\right]D_{F}({x}_{k},{\mathsf{x}}_{*})$
@@
We know that $2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})=\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}]$. From Lemma B.7, Equation (2.1), and Equation (5.3), it follows that
@@
$=\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}],\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])\,|\,\mathfrak{F}_{k}]$
@@
$\leq 2\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-\mathsf{E}_{k}[D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])].$
@@
The last inequality comes from Lemma B.5. ∎
Proof of Lemma 5.14. From Lemma B.6, we know that $D_{h}({x}_{k},{z}_{k+1})\leq(S_{1}+S_{2})/2$, where
@@
Using the gain function in Equation (2.1), then Lemma B.5, and finally Equation (5.7), we have
@@
$\leq 4L\alpha_{k}^{2}\mathsf{E}_{k}\left[G\left(\nabla h({x}_{k}),\nabla h({x}_{k}),\frac{1}{L}(\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*}))\right)\right]D_{F}({x}_{k},{\mathsf{x}}_{*})$
@@
We know that $2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})=\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}]$. From Lemma B.7, Equation (2.1), and Equation (5.8), it follows that
@@
$=\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}],\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])\,|\,\mathfrak{F}_{k}]$
@@
$\leq 2\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-\mathsf{E}_{k}[D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])].$
@@
The last inequality comes from Lemma B.5. ∎
Proof of Lemma A.2. Thanks to Lemma B.6, we know that $D_{h}({x}_{k},{z}_{k+1})\leq(S_{1}+S_{2})/2$, where
@@
Using the gain function in Equation (2.1), then Lemma B.5, and finally Equation (A.1), we have
@@
$\leq 4L\alpha_{k}^{2}\mathsf{E}\left[G\left(\nabla h({x}_{k}),\nabla h({x}_{k}),\frac{1}{L}(\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*}))\right)\,|\,\mathfrak{F}_{s,k}\right]D_{F}({x}_{k},{\mathsf{x}}_{*})$
$\leq 4L\alpha_{k}^{2}G_{k}D_{F}({x}_{k},{\mathsf{x}}_{*}).$
We know that $2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})=\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}]$. From Lemma B.7, Equation (2.1), and Equation (A.2), it follows that
$\mathsf{E}[S_{2}\,|\,\mathfrak{F}_{s,k}]$ $=\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}],\nabla h({x}_{k}))\,|\,\mathfrak{F}_{s,k}]$
$=\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}],\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])]$
$\qquad-\mathsf{E}_{k}[D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])\,|\,\mathfrak{F}_{s,k}]$
@@
$\qquad-\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}])\,|\,\mathfrak{F}_{s,k}]$
$\leq 2\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}])\,|\,\mathfrak{F}_{s,k}].$
@@
Proof of Theorem A.5. Proposition B.1 and Lemma B.7 give
$\mathsf{E}[D_{h}({\mathsf{x}}_{*},{x}_{k+1})]$ $\leq D_{h}({\mathsf{x}}_{*},{x}_{k})+\alpha_{k}^{2}B_{k}\sigma_{k}^{2}$
$\qquad-\alpha_{k}\left[1-\alpha_{k}A_{k}\right]\mathsf{E}[F({x}_{k})-F({\mathsf{x}}_{*})]$
@@
$\qquad+2L\alpha_{k}^{2}B_{k}\mathsf{E}[D_{F}(\tilde{{x}}_{s},{\mathsf{x}}_{*})].$ (B.16)
In the last inequality, we used Lemma B.5. By taking total expectation, it follows
@@
Then by summing over $k=0,\ldots,m-1$, taking the total expectation and recalling Remark A.4, we obtain
@@
It still holds by choosing $\tilde{{x}}_{s+1}=\sum_{k=0}^{m-1}\frac{1}{m}{x}_{k}$ , in Algorithm A.1, and using Jensen inequality to lower bound $\sum_{k=0}^{m-1}F({x}_{k})$ by $mF(\tilde{{x}}_{s+1})$. We finally have

### v2

@@ B.2 Technical lemmas
Suppose that Assumptions 4.5() are verified and that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1() with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence. Then, for all $k\in\mathbb{N}$,
@@
Define $\mathsf{E}_{k}[\cdot]=\mathsf{E}[\cdot\,|\,\mathfrak{F}_{k}]$. It follows from Proposition B.1() that
@@
$=\frac{1}{2}\Big[D_{h^{*}}\left(\nabla h({\mathsf{x}}_{1}^{+}),\nabla h({\mathsf{x}})\right)+D_{h^{*}}\left(\nabla h({\mathsf{x}}_{2}^{+}),\nabla h({\mathsf{x}})\right)\Big].$
@@
Lemma B.7 (Bregman variance decomposition by [24, Theorem 2.3]).
Let x{x} be a random variable on $\operatorname*{dom}h^{*}$ such that $\mathsf{E}[{x}]\in\operatorname*{int}\operatorname*{dom}h^{*}$. Then for any ${\mathsf{u}}\in\operatorname*{int}\operatorname*{dom}h^{*}$,
@@
B.3 Proofs of Section 4.1()
Proof of Theorem 4.6(). From Lemma B.4() and since $B_{k}-\rho M_{k}\leq 0$ , we get
@@
Proof of Theorem 4.7(). From Lemma B.4() and relative strong convexity of $F$, we have
@@
B.4 Proofs of Section 5()
Proof of Theorem 5.2(). From Proposition B.1(), we have
@@ B.4 Proofs of Section 5()
Proof of Theorem 5.4(). Proposition B.1() gives
@@
Proof of Lemma 5.8(). From Equation (4.2), we infer that $\nabla h({z}_{k+1})=\nabla h({x}_{k})-\alpha_{k}\nabla f_{i_{k}}({x}_{k})+\alpha_{k}{e}_{k}$. Set $\nabla h({\mathsf{x}}_{1}^{+})=\nabla h({x}_{k})-2\alpha_{k}[\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*})]$ and $\nabla h({\mathsf{x}}_{2}^{+})=\nabla h({x}_{k})-2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})$. Then, from Lemma B.6(), we know that $D_{h}({x}_{k},{z}_{k+1})\leq(S_{1}+S_{2})/2$, where
@@
Using the gain function in Equation (2.1), then Lemma B.5(), and finally Equation (5.2), we have
@@
$\leq 4L\alpha_{k}^{2}G_{k}\mathsf{E}_{k}\left[D_{h^{*}}\left(\nabla h({x}_{k})-\frac{1}{L}(\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*})),\nabla h({x}_{k})\right)\right]$
@@
We know that $2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})=\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}]$. From Lemma B.7(), Equation (2.1), and Equation (5.3), it follows that
@@
$=\mathsf{E}_{k}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}],\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])]$
$\qquad-D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])$
@@
$\leq 2\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}]).$
@@
The last inequality comes from Lemma B.5(). ∎
Proof of Lemma 5.15(). From Lemma B.6(), we know that $D_{h}({x}_{k},{z}_{k+1})\leq(S_{1}+S_{2})/2$, where
@@
Using the gain function in Equation (2.1), then Lemma B.5(), and finally Equation (5.7), we have
@@
$\leq 4L\alpha_{k}^{2}G_{k}\mathsf{E}_{k}\left[D_{h^{*}}\left(\nabla h({x}_{k})-\frac{1}{L}(\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*})),\nabla h({x}_{k})\right)\right]$
@@
We know that $2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})=\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}]$. From Lemma B.7(), Equation (2.1), and Equation (5.8), it follows that
@@
$=\mathsf{E}_{k}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}_{k}[\zeta_{k}],\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])]$
$\qquad-D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])$
@@
$\leq 2\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}]).$
@@
The last inequality comes from Lemma B.5(). ∎
Proof of Lemma A.2(). Thanks to Lemma B.6(), we know that $D_{h}({x}_{k},{z}_{k+1})\leq(S_{1}+S_{2})/2$, where
@@
Using the gain function in Equation (2.1), then Lemma B.5(), and finally Equation (A.1), we have
@@
$\leq 4L\alpha_{k}^{2}G_{sm+k}\mathsf{E}\left[D_{h^{*}}\left(\nabla h({x}_{k})-\frac{1}{L}(\nabla f_{i_{k}}({x}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*})),\nabla h({x}_{k})\right)\,|\,\mathfrak{F}_{s,k}\right]$
$\leq 4L\alpha_{k}^{2}G_{sm+k}D_{F}({x}_{k},{\mathsf{x}}_{*}).$
We know that $2\alpha_{k}(\nabla f_{i_{k}}({\mathsf{x}}_{*})-{e}_{k})=\zeta_{k}-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}]$. From Lemma B.7(), Equation (2.1), and Equation (A.2), it follows that
$\mathsf{E}[S_{2}\,|\,\mathfrak{F}_{s,k}]$ $=\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}],\nabla h({x}_{k}))\,|\,\mathfrak{F}_{s,k}]$
$=\mathsf{E}[D_{h^{*}}(\nabla h({x}_{k})+\zeta_{k}-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}],\nabla h({x}_{k})-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}])\,|\,\mathfrak{F}_{s,k}]$
$\qquad-D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}])$
@@
$\qquad-D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}])$
$\leq 2\alpha_{k}^{2}G_{sm+k}\sigma_{k}^{2}-D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}]).$
@@
Proof of Theorem A.4(). Proposition B.1() and Lemma B.7() give
$\mathsf{E}[D_{h}({\mathsf{x}}_{*},{x}_{k+1})\,|\,\mathfrak{F}_{s,k}]$ $\leq D_{h}({\mathsf{x}}_{*},{x}_{k})+\alpha_{k}^{2}B_{k}\sigma_{k}^{2}$
$\qquad-\alpha_{k}\left[1-\alpha_{k}A_{k}\right]\mathsf{E}[F({x}_{k})-F({\mathsf{x}}_{*})\,|\,\mathfrak{F}_{s,k}]$
@@
$\qquad+2L\alpha_{k}^{2}B_{k}D_{F}(\tilde{{x}}_{s},{\mathsf{x}}_{*}).$ (B.16)
In the last inequality, we used Lemma B.5(). By taking total expectation, it follows
@@
Then by summing over $k=0,\ldots,m-1$ and recalling Remark A.3(), we obtain
@@
It still holds by choosing $\tilde{{x}}_{s+1}=\sum_{k=0}^{m-1}\frac{1}{m}{x}_{k}$ , in Algorithm A.1(), and using Jensen inequality to lower bound $\sum_{k=0}^{m-1}F({x}_{k})$ by $mF(\tilde{{x}}_{s+1})$. We finally have
@@ Appendix C Full results of the numerical experiments
Figure 4: The case of diagonal AA: closed-form solution to the proximal mapping. There is no instability for BSPPA in that case, showing that its instability in Figure 1() is due to the inexact proximal mapping. This proves it is less stable to inexact algorithms compare to its variance-reduced counterparts.